Teenage suicide is a serious problem in the United States. The family economic status of a sample of 38 teenage who committed suicide was distributed as follows:

ClassUpperUpper MiddleLower MiddleLowerNumber of Suicides141086

Is there statistical evidence that the proportions in the classes differ at alpha 0.05?

  1. Formulate the H0 and H1
  2. C.V. at alpha 0.05 level of significance
  3. Compute the test value: Chi-square X2
  4. Decision
  5. Interpretation
1
Expert's answer
2022-01-14T07:08:33-0500

1.

"H_0:" the proportions in the classes do not differ

"H_a:" the proportions in the classes differ


2.

(Rows-1)*(Columns-1) = (2-1)*(2-1) = 1

critical value:

"\chi^2_{crit}=3.84"


3.

"\chi^2=\sum \frac{(O_i-E_i)^2}{E_i}=0.005"


where "O_i" is observed frequency,

"E_i" = (row total)(column total)/(total) is expected frequency


4.

Since "\chi^2<\chi^2_{crit}" we accept null hypothesis


5.

The proportions in the classes do not differ.